OpenStudy (anonymous):

Solve using the substitution method x + 3y = −8, 4x − 3y = 23

OpenStudy (anonymous):

x=3 y=-11/3

OpenStudy (anonymous):

ho u got that

OpenStudy (anonymous):

how

OpenStudy (angela210793):

x + 3y = −8 add both sides......5x=15--->x=3--->3+3y=-8 4x − 3y = 23 3y=-11 y=-11/3

OpenStudy (anonymous):

3y= -8-x sub that into the second equation to get 4x - (-8-x) = 23 open brackets to get 5x = 15 x = 3 then 3 + 3y = -8 3y = -11 y = -11/3

OpenStudy (angela210793):

hmm ok....but the 1st solution is kinda easier cause -3y and 3y r simplified

OpenStudy (anonymous):

k